OpenStudy (anonymous):

Determine whether the function \[f(x)=\sin^{-1} x\] satisfies the conditions of the Mean Value Theorem on the interval [0,1/2]. If so find the points guaranteed to exist by the theorem

OpenStudy (zehanz):

Here is a graph of the situation (in red). The dotted line is the line between the endpoints. Conditions of the MVT: 1. contiuity on [0, 1/2]: OK 2. differentiability on (0, 1/2): OK So the conditions are met. Now look for a point on the graph where the tangent line has the same slope as the dotted line.

OpenStudy (anonymous):

that is a hard look. you need to solve \[f'(x)=\frac{f(b)-f(a)}{b-a}\] for \(x\)

OpenStudy (anonymous):

the right hand side is \(\frac{\sin^{-1}(\frac{1}{2})-\sin^{-1}(0)}{\frac{1}{2}-0}\)

OpenStudy (anonymous):

once you evaluate, you should get \[\frac{\pi}{6}\times 2=\frac{\pi}{3}\]

OpenStudy (anonymous):

oh I see thanks guys

OpenStudy (anonymous):

you still have to take the derivative, set it equal to \(\frac{\pi}{3}\) and solve

OpenStudy (anonymous):

yea I know that but i got the wrong answer for the first part now that I see where the mistake was made the next part is easy \[\sqrt{1-\frac{ 9 }{ \pi ^{2} }}\]

OpenStudy (zehanz):

@satellite73: Although the graph of \(sin^{-1}x\) is almost straight, the dotted line is still visible and helps solving the problem imo :)